Solve the equation for x.
2(x-5)^2=3

Answer by checkley75(3666)

2(x-5)^2=3
2(x^2-10x+25)=3
2x^2-20x+50-3=0
2x^2-20x+47=0using the quadratic equation we get:
x=(20+-sqrt[-20^2-4*2*47])/2*2
x=(20+-sqrt[400-376])/4
x=(20+-sqrt24)/4
x=(20+-4.9)/4
x=(24.9)/4
x=6.22 answer.
x=(20-4.9)/4
x=15.1/4
x=3.78 answer.